INFRARED EMISSION SPECTRUM OF DBO

\underline{PINA COLARUSSO}, KEQING ZHANG, PETER BERNATH, Department of Chemistry, University of Waterloo, Waterloo, Ontario, Canada N2L 3G1; YOSHIYUKI KAWASHIMA, Department of Chemical Technology, Kanagawa Institute of Technology, Atsugi, Kanagawa 243-02, Japan; and EIZI HIROTA, The Graduate University for Advanced Studies, Hayama, Kanagawa 240-01, Japan.

We generated the transient molecule DBO by reacting deuterium with the parent molecule, HBO, at temperatures above 800^{∘}C in a ceramic tube furnace. Infrared emission spectra were recorded with a Fourier transform spectrometer at a resolution of 0.01 cm^{−1} from 350 to 3600 cm^{−1}. We observed the three fundamental vibrational bands as well as various hot and combination bands. A strong Coriolis interaction was detected between the ν_{1} and ν_{2}+ν_{3} energy levels. The high-resolution infrared data for DBO will be combined with previous results^{a}^{,}^{b} to yield a thorough analysis of the ground electronic state.

aY. Kawashima et al., J. Molec. Spectrosc. 133, 116-127 (1989) and 44^{th} Symp. Molec. Spectrosc., Columbus, (1989).

bP. Colarusso et al., 51^{st} Sympos. on Molec. Spectrosc., Columbus, (1996).